International Candy Inc. operated at 100% of capacity during its first month and incurred the following costs:

Production costs (10,000 units):

Direct materials

$140,000

Direct labor

40,000

Variable factory overhead

20,000

Fixed factory overhead

4,000

$204,000

Operating expenses:

Variable operating expenses

$ 34,000

Fixed operating expenses

2,000

36,000

If 2,000 units remain unsold at the end of the month and sales total $300,000 for the month, what would be the amount of income from operations reported on the variable costing income statement?

(Each answer should have possibilities for answers entered with or without dollar signs and commas. Thus the correct answer for a problem could be $100,000 100,000 100000 $10000)

In order to earn credit for your answers to the problems, you are required to show all supporting work.
Answer #1

Amount of income from operations reported on the variable costing income statement = Sales - Variable Costs - Fixed Costs

= 300000 - (8000 x (14+4+2+3.4)) - (4000 + 2000)

= 300000 - (8000 x 23.4)) - 6000

= 300000 - 187200 - 6000

= $106,800

Note: Variable Expenses per unit = Variable Expense/10000 units

Variable costs are calculated for 8000 units only as they are sold but fixed expenses are expensed as incurred in variable costing.